Please solve this inequasion for me!!! |x+2|/|x-3|≤2
Answer by stanbon(75887)

|x+2|/|x-3|≤2
------------------
1st: x cannot be 3
----------------------------------
2nd: Solve the EQUALITY:
(x+2)/x-3) = 2
x+2 = 2x-6
x = 8
-------
If -2 < x < 3,
(x+2)/[-(x-3)] = 2
x+2 = -2x+6
3x = 4
x = 1.3333333...
-------------
Plot those three values on a number line: 1.3333, 3, 8
------------------
Check test values in the 4 resulting intervals to see where the solutions are.
-------------
Check x = -10: |-10+2|/|-10-3| < = 2; true, so solutions in (-inf,1.3333]
Check x = 2 ; |2+2|/|2-3| = 4 < 2; false
Check x =5 ; |5+2|/|5-3| =7/2 < 2; false
Check x = 10 ; |10+2|/|10-3|=12/7 < 2; true, so solution in [8,+inf)
--------------
Ans: (-inf,1.333]U[8,+inf)
==============================
Cheers,
Stan H.